OpenStudy (anonymous):

Dawn deposits $850.00 in an account that earns 6% simple interest. About how many years will it take for her account balance to earn $250.00 in interest?

OpenStudy (johnweldon1993):

Simple interest uses the formula I = PRT I = interest earned (250.00) P = principal/original rate (850.00) R = interest rate (.06) T = time (in years) (what we want) so...lets rearrange this equation for T since that is what we want \[I = prt\] Divide both sides by 'pr'...this will isolate 't' and allow us to solve for it \[\frac{ I }{ pr } = \frac{ prt }{ pr }\] This leaves us with \[t = \frac{ I }{ pr }\] So now input your values for I, P, R and solve for 't'...

OpenStudy (anonymous):

would the answer be 5 years?

OpenStudy (johnweldon1993):

Since the question says "about how many years" yes the answer is 5 years...good job!!

OpenStudy (anonymous):

Thank you I have another question could you help me.?

OpenStudy (johnweldon1993):

If I can sure...

OpenStudy (anonymous):

Joe deposits $50,000.00 in an account that earns 8.3% simple interest. About how many years will it take for his account balance to earn $80,000.00 in interest?

OpenStudy (johnweldon1993):

So it's basically the same as what we just did....can you attempt it first and see if you can come to an answer?

OpenStudy (anonymous):

I tried 50,00000 dived by 415000 and it made no sense to me

OpenStudy (johnweldon1993):

Remember when you simplify the formula it comes to \[t = \frac{ I }{ pr }\] I = 80000 P = 50000 R = .083

OpenStudy (anonymous):

I have that correct, but I am stumped now

OpenStudy (johnweldon1993):

Why are you stumped? Ahh I see what you did....remember that the rate is 8.3% ....you don't multiply by this....you need to change this into a decimal \[\frac{ 8.3 }{ 100 } = .083\] What you did was multiply 50000 by 8.3... but it should be 50000 by .083

OpenStudy (anonymous):

Then I multiplied 50000.x 0.83 = 4150;

OpenStudy (johnweldon1993):

I'm assuming you meant .083...but yes that is correct....so now you have \[t = \frac{ 80000 }{ 4150 }\] so what is t...

OpenStudy (anonymous):

There is my problem, I placed 50000. instead of the 80000.

OpenStudy (johnweldon1993):

Ahh....yes remember I = 80000 and P = 50000 so now you're all good....t = ?

OpenStudy (anonymous):

answer is 19.years

OpenStudy (johnweldon1993):

That would be correct

OpenStudy (anonymous):

Thank you so much , that was driving me crazy, You have a great 4th.

OpenStudy (johnweldon1993):

No problem! and you as well!!